### v3

@@ Abstract
Recent work demonstrates that, after being fine-tuned on a high-quality instruction dataset, the resulting model can obtain impressive capabilities to address a wide range of tasks. However, existing methods for instruction data generation often produce duplicate data and are not controllable enough on data quality. In this paper, we extend the generalization of instruction tuning by classifying the instruction data to 4 code-related tasks which is specifically designed for enhancing instruction tuning of Code Large Language Models (LLMs) and propose a LLM-based Generator-Discriminator data process framework to generate diverse, high-quality instruction data from open source code. Hence, we introduce CodeOcean, a dataset comprising 20,000 instruction instances across 4 universal code-related tasks, which is aimed at augmenting the effectiveness of instruction tuning and improving the generalization ability of fine-tuned model. Subsequently, we present WaveCoder, a fine-tuned Code LLM with Widespread And Versatile Enhanced instruction tuning. Our experiments demonstrate that WaveCoder outperforms other open-source models in terms of generalization ability across different code-related tasks at the same level of fine-tuning scale. Moreover, WaveCoder exhibits high efficiency in previous code generation tasks. This paper thus offers a significant contribution to the field of instruction data generation and fine-tuning models, providing new insights and tools for enhancing performance in code-related tasks.
@@ 1 Introduction
In recent months, numerous Large Language Models (LLMs) such as ChatGPT, GPT-4 [1], and Gemini 111https://deepmind.google/technologies/gemini have attained unprecedented performance levels in a broad array of NLP tasks. These models utilize a self-supervised pre-training process, and subsequent fine-tuning, to demonstrate exceptional zero/few-shot capabilities, effectively following human instructions across an array of diverse tasks.
However, the overhead of holding such a large model is huge. Therefore, some relatively small LLM, especially Code Large Language Models (Code LLMs), have attracted considerable attention due to their exceptional performance on a wide range of code-related tasks. Given that LLMs can accrue a wealth of professional knowledge through pre-training [1, 2, 3], an efficient and effective pre-training process on a code corpus is critical for Code LLMs. Several previous works, including Codex [4], CodeGen [5], StarCoder [6], and CodeLLaMa [7], have successfully demonstrated that this process can significantly improve their capability to tackle code-related problems.
Additionally, instruction tuning has been shown to perform at a satisfactory level across various tasks, as demonstrated in research based on instruction-tuning like FLAN [8], ExT5 [9], and FLANT5 [10]. These studies incorporated thousands of tasks into their training pipelines to improve the generalization ability of pre-trained model on downstream tasks. InstructGPT [11] aligns user input effectively by integrating high-quality instruction data written by human labelers, which inspires further exploration into the utility of instruction tuning. Stanford Alpaca [12] utilized ChatGPT to generate instruction data via a self-instruct method, which in turn was used in the instruction tuning process. Further improvements to the effectiveness of instruction tuning have been explored in WizardLM [13] and WizardCoder [14], where the evol-instruct method was applied. These works collectively exemplify the potential of instruction tuning in enhancing the performance of large language models.
Based on these work, our intuition is that instruction-tuning can activate the potential of LLMs then align the pre-trained models to an outstanding intelligent level. Therefore, we have summarized the main function of instruction tuning:
Generalization. Initially proposed to augment the cross-task generalization capabilities of Language Models (LMs) [8, 9, 10], instruction-tuning enhances performance across a multitude of unseen tasks when fine-tuned with diverse NLP task instructions.
Alignment. Pre-trained models [2, 3, 15, 16], having learned from a vast number of self-supervised tasks at both token and sentence levels, already possess the capability to comprehend text inputs. Instruction-tuning offers instruction-level tasks for these pre-trained models[15, 17], enabling them to extract more information from the instruction beyond the semantics of raw text. This additional information is user intention and enhances their interactive abilities with human users, thereby contributing to alignment.
To enhance the performance of Code LLMs through instruction tuning, many existing methods for generating instructional data have been devised, focusing primarily on the two aforementioned aspects. For example, self-instruct [12], evol-instruct [13] employed the zero/few-shot ability of teacher LLM to generated instruction data, which present an amazing way for instruction data generation. However, these generation ways are too dependent on the performance of the teacher LLM and sometimes produces a large amount of duplicate data which will reduce the efficiency of fine-tuning [18, 19, 20]. Octopack [21] constructs a code instruction dataset leveraging the natural structure of Git commits to break away from dependence on teacher LLMs. Howerver, Ensuring the quality of data in git messages presents a considerable challenge, and the comprehensive screening of data through artificial filtering rules is often a complex task.
To solve these problems, as shown in Figure 2, we propose a method that could make full use of source code and explicitly control the quality of generated data. Owing to the fact that instruction tuning is to align the pre-training model to the instruction-follow training set [1, 17], we prensent a LLM Generator-Disciminator framework for instruction data generation. By employing generation and discrimination, our method can make the data generation process more customizable and more controllable. Taking the raw code as input and selecting the core dataset, our method could stably generate more realistic instruction data and control the diversity of data by adjusting the distribution of the raw code.
@@
In this paper, our focus is on enhancing the performance of code LLMs by applying instruction-tuning. Addressing the aforementioned challenges, we refine the instruction data by classifying the instruction instances to four universal code-related tasks: Code Summarization, Code Generation, Code Translation, and Code Repair and generate a dataset of 20,000 instruction instances, termed CodeOcean, for the four code-related tasks with our data generation strategy.
To validate our approach, taking StarCoder [6], CodeLLaMa [7], and DeepseekCoder [22] as our base models, we introduce WaveCoder models, fine-tuned with our data generation strategy and evaluate them on HumanEval [4], MBPP [23], HumanEvalPack [24] benchmarks, experimental results show that our WaveCoder exhibits exceptional performance based on a small-scale instruction tuning.
2 CodeOcean: Four-task Code-related Instruction Data
In this section, we introduce the methodological details of our exploration. We firstly choose 4 representative coding tasks and collect raw code from open source dataset. For each task, we employ GPT-3.5-turbo to generate instruction data for fine-tuning. The generation prompt is shown in Table 2 and the data generation process is shown in Figure 3.
2.1 Tasks Setting
@@
Given the breadth of code-related tasks from CodeXGLUE [25], we have selected four of the most universally representative and common tasks from the three generative tasks (code-to-text, text-to-code, and code-to-code) for further exploration. These include Code Summarization, Code Generation, Code Translation, and Code Repair. Detailed descriptions of these tasks can be found below.
@@
Code Summarization (code-to-text). This task aims to create a brief summary of a given code. The raw code is used as input and the model’s response is reformulated into an instructional format.
Code Generation (text-to-code, code-to-code). In this task, the model is expected to generate code based on a user’s demand description. The model generates instructions to simulate the user’s demand discourse based on the raw code. The generated raw code is then considered as the output.
Code Translation (code-to-code). This task involves converting one programming language into another. A prompt and raw code are given to the model, and it generates instructions and the translated code.
Code Repair (code-to-code). The aim of this task is to provide correct code based on potential issues in the given code. The model is expected to generate solutions for the raw code, typically within one or two sentences, which are then taken as the output.
2.2 Generation of Training Data
In past research work [26, 27], many researchers have discovered that data quality and diversity play an important role in the instruction tuning process rather than data amount. The improvement of data quality and diversity are directly related to the performance of the fine-tuned model. Therefore, to ensure the data quality and diversity of instruction instance, we propose: 1) a method that can retain the diversity of raw code to the utmost extent. 2) a LLM-based Generatior-Discriminator framework to further improve the data quality of instruction instances.
@@ 2.2.1 Raw Code Collection
To ensure the quality and diversity of raw code, we manually define some filtering rules and utilize a cluster method KCenterGreedy [28, 29] to get the raw code collection from the open source code dataset for instruction data generation. In this work, we select CodeSearchNet 222https://huggingface.co/datasets/code_search_net which contains 2 million (comment, code) pairs from open-source libraries hosted on GitHub and includes code and documentation for 6 programming languages as our foundation dataset and develop a series of manually filtering rules to ensure the basic data quality of raw code. Subsequently, we apply the coreset-based selection method KCenterGreedy to maximize the diversity of the raw code. More details are as follows:
Manually defined filtering rules. In order to select high-quality code for instruction-tuning, we make some rules to filter the foundation dataset preliminarily. The main rules are as follows:
i)
The length of required code should not be too long or too short. In this work, we filtered the code whose length lower than 50 or higher than 800.
ii)
Followed Code Alpaca333https://github.com/sahil280114/codealpaca, we have eliminated the raw code containing words from the blacklist, since these code could potentially reduce the performance of the resulting model.
Coreset-based selection method. After filtered by the above rules, the size of the raw code dataset dropped to 1.2 million, which is still a huge scale and too expensive to fine-tune. To this end, we employed KCenterGreedy algorithm based on code embeddings to reduce the amount of training data while maximizing data diversity of raw code as much as possible.
In this method, our main work is divided into two steps:
i)
Code embedding. In this step, we encode all the raw code samples using a pre-trained language model (roberta-large-v1 [16] in this paper). we take the [CLS] token as the code embedding for one input raw code, which is not defined in generative language models like GPT-3.5.
ii)
Coreset sampling. In the coreset sampling step, our goal is to use a small number of samples to represent the distribution of the overall data, thus we employed unsupervised clustering and focused on the data representations. After obtaining the code embeddings from the first step, we used KCentergreedy algorithm [28] to select the core dataset and maximize data diversity, which aims to choose center points such that minimize the largest distance between a random data point and its nearest center, and has been proven efficient in obtaining a set of core samples of one distribution.
Task | Num | Per(%) | Prompt |
Code Generation | 11370 | 57.0 | Implementing functions that perform specific operations given input. |
Code Summarization | 3175 | 15.9 | Write clear and concise documentation for the given code. |
Code Repair | 3144 | 15.7 | Identify and fix errors in the given code. |
2.2.2 LLM-based Generator-Discriminator Framework
After the process of collecting raw code, it is more significant to generate high-quality and diverse instruction data and reformat them to training data. To this end, shown in Figure 2, we propose a novel LLM-based Generator-Discriminator Framework, where the Generator can leverage an extensive amount of unsupervised open source code to generate supervised instruction data. and the Discriminator can continue analyzing and filtering the instruction data. By incorporating such a framework into the open source code, the diversity of the generated data no longer relies solely on capability of the teacher LLM itself and the data quality of instruction data generation can be enhanced by the Discriminator.
Generation Phase. In the generation phase, we utilize GPT-4 to generate a task definition in the context of a specific scenario. As shown in Figure 3, following the model-generated task definition, we manually develop the generation requirements for the each code-related task. Integrating both the task definition and all associated requirements into the generation prompt, we take the raw code as an input and select good and bad case examples from the example database to leverage the few-shot abilities of GPT-3.5 to generate the necessary knowledge and information for instruction tuning. However, based on our experiments, we observe that few-shot generation is not always the best way to get instruction data especially on code generation task. We analyzed that this is because the teacher LLM always imitated the given example during the few-shot generation process, which will damage the diversity of instruction data. Therefore, we employed zero-shot generation on code generation task and few-shot generation on other tasks.
Discrimination Phase. During the exploration of the instruction generation process, we noticed that the data quality of the instruction instances cannot be guaranteed through the generation phase alone. In order to enhance the controllability of data generation and further ensure the data quality, we designed a LLM-based discriminator to continue analyzing and filtering the instruction data. For the discrimination phase, we establish a series of criteria related to instruction data and employ GPT-4 to assess the quality of instruction instances. However, due to the large number of rules, the discriminator cannot conduct effective and comprehensive discrimination. To solve this problem, we disassemble the discrimination rules to some subtopics inspired by Zero-shot-CoT [30] and employ the LLM-based discriminator to analyze the generation step by step . By adopting this method, the discrimination process can be more contrallable when filtering low-quality instruction instances. After the discrimination process, each instruction instance is classified as either a good or bad case and the classification information is subsequently reused in the following generation as examples. For the reusage of these classified instruction instance, different from self-instruct [31] which solely utilize the exemplification as good example, we exploit both the good generation and bad generation as the few-shot example so that the generator can gain from the analysis of bad example. An example of our Generator and Discriminator framework is presented in Figure 4. This framework provides a comprehensive approach to generating and evaluating instruction data, ensuring a high-quality and diverse training dataset.
3 Experiments
3.1 Training
Unlike the previous work [14, 32, 33] that mainly focus on code generation task, we hope that CodeOcean could gain more general problem-solving capabilities on a wider range of coding tasks. Therefore, our training relies on a 20k instruction dataset generated by our LLM-based Generator-Discriminator framework and an additional logical dataset. The 20k dataset covers 4 common code tasks, including code generation, code translation, code summarization, and code repair.
To obtain WaveCoder models, We choose StarCoder-15B, CodeLLaMa (7B and 13B), DeepseekCoder-6.7B as the base model and fine-tune all the base model for 3 epochs using NVIDIA A100-80GB GPU. For StarCoder-15B and CodeLLaMa-13B, we set the global batch size to 128 using Tensor Parallel and set the initial learning rate at 2e-5. For DeepseekCoder-6.7B and CodeLLaMa-7B, we set the global batch size to 256 using the Fully Sharded Data Parallel (FSDP) module from Pytorch and set the initial learning rate at 5e-5. The sequence length of our each instruction sample is less than 1024. Followed Alpaca [12], we set the fine-tuning prompt as follows:
3.2 Evaluation
@@
OctoCoder [21] | 15B | StarCoder | 13K | 46.2 | 43.5 |
WizardCoder [14] | 15B | StarCoder | 78K | 57.3 | 51.8 |
WaveCoder-SC-15B ![]() |
@@
CodeLLaMa [7] | 7B | - | ✘ | 33.5 | 41.4 |
CodeLLaMa-instruct [7] | 7B | CodeLLaMa | 14K | 34.8 | 44.4 |
WaveCoder-CL-7B ![]() |
@@
CodeLLaMa [7] | 13B | - | ✘ | 36.0 | 47.0 |
CodeLLaMa-instruct [7] | 13B | CodeLLaMa | 14K | 42.5 | 49.4 |
WaveCoder-CL-13B ![]() |
@@
DeepseekCoder [22] | 6.7B | - | ✘ | 49.4 | 60.6 |
Magicoder-DS[34] | 6.7B | DeepseekCoder | 75K | 66.5 | 60.4 |
WaveCoder-DS-6.7B ![]() |
@@
3.2.1 Benchmarks and Baselines
Proprietary Models. We present the self-reported results from an array of SoTA LLMs, including ChatGPT (gpt-3.5-turbo), GPT-4. If not reported, we use the results from Octopack [21] or evaluate by ourselves.
Open Source Models. Base models comprise StarCoder, CodeLLaMa and Deepseek-Coder with prompting. To ensure an equitable comparison, we opted to select those instructed models that have been trained using fewer than 100,000 instruction instances for our comparative analysis.
3.2.2 metrics
Pass@k. Followed most previous works, we also adopt the pass@k metric proposed by OpenAI [4] to evaluate the performance of our model. By generating samples for each promblem and checking the correctness for each sample with the test cases, the pass@k score can be calculated by the unbiased estimator:
where is the total number of sample generated for a problem, is the number of correct samples passing all the test cases.
3.2.3 Evaluation on Code Generation Task
HumanEval and MBPP are two representative benchmarks for code generation task where the model should generate the complete code based on the function signature and the docstring of the problem. Table 3 shows the pass@1 score of different LLMs on both benchmarks. From the results, We have the following observations:
WaveCoder models substantially outperform the instructed models training with fewer than 20K instruction tuning data (InsT Data) and CodeAlpaca with 20K data. Following the fine-tuning process, the performance of our models exhibit substantial improvement when compared to both the foundation model and a selection of open-source models, but it still lags behind proprietary models the instructed models training with more than 70K training data.
Refined and diverse instruction data can significantly improve the efficiency of instruction tuning. According to the comparison in Table 3, WaveCoder solely relies on a dataset encompassing 20,000 entries to accomplish performance levels on par with those of Magicoder [34] and Wizardcoder [14]. Although WaveCoder falls behind WizardCoder [14] and Magicoder [34] on code generation task, there is a considerable difference in data size among them. Nevertheless, the performance lag of WaveCoder is relatively minor.

### v4

@@ Abstract
Recent work demonstrates that, after instruction tuning, Code Large Language Models (Code LLMs) can obtain impressive capabilities to address a wide range of code-related tasks. However, current instruction tuning methods for Code LLMs mainly focus on the traditional code generation task, resulting in poor performance in complex multi-task scenarios. In this paper, we concentrate on multiple code-related tasks and present WaveCoder, a series of Code LLMs trained with Widespread And Versatile Enhanced instruction data. To enable the models to tackle complex code-related tasks, we propose a method to stably generate diverse, high-quality instruction data from open source code dataset in multi-task scenarios and obtain CodeSeaXDataset, a dataset comprising 19,915 instruction instances across 4 code-related tasks, which is aimed at improving the generalization ability of Code LLM. Our experiments demonstrate that WaveCoder models significantly outperform other open-source models in terms of the generalization ability across different code-related tasks. Moreover, WaveCoder-Ultra-6.7B presents the state-of-the-art generalization abilities on a wide range of code-related tasks.
@@ 1 Introduction
Recently, Large Language Models (LLMs) such as ChatGPT, GPT-4 OpenAI (2023), and Gemini 111https://deepmind.google/technologies/gemini have attained unprecedented performance levels in a broad array of NLP tasks. These models utilize a self-supervised pre-training process, and subsequent supervised fine-tuning to demonstrate exceptional zero/few-shot capabilities, effectively following human instructions across various tasks.
For code-related tasks, several previous works, including Codex Chen et al. (2021), StarCoder Li et al. (2023a), CodeLLaMa Roziere et al. (2023) and DeepseekCoder Guo et al. (2024), have successfully demonstrated that pre-training on code corpus can significantly improve the model’s capability to tackle code-related problems. After the process of pre-training, instruction tuning Wei et al. (2022); Aribandi et al. (2022); Chung et al. (2022) has shown its effectiveness in the aspect of improving the quality of LLM responses. To specifically enhance the performance of Code LLMs on code-related tasks through instruction tuning, many existing methods for instruction data generation have been designed. For example, Code Alpaca Chaudhary (2023) utilizes the method of self-instruct Wang et al. (2023a) within the coding domain, leveraging the few-shot capabilities of teacher LLM to generate instruction data. Similarly, WizardCoder Luo et al. (2024) applies the evol-instruct Xu et al. (2024) approach based on Code Alpaca, demonstrating a novel and effective method for the generation of instruction data. These applications underscore the potential of utilizing teacher LLMs to produce instructional content effectively, thereby offering an avenue for the creation of instruction data in the code domain. However, the quality of the data they generate heavily relies on the performance of the teacher LLM and the limited initial seeds, which often produces a large amount of duplicate instruction instances and reduce the effectiveness of instruction tuning Xu et al. (2022); Yan et al. (2023); Lee et al. (2022). To break away from dependence on teacher LLMs, Octopack Muennighoff et al. (2024) constructs a code instruction dataset leveraging the natural structure of Git commits. Nonetheless, ensuring the quality of data in git messages presents a considerable challenge, and the comprehensive screening of data through artificial filtering rules is often a complex task. Additionally, these endeavors are predominantly centered on traditional code generation tasks and lack the capability to produce detailed, task-specific instructions in multi-task scenarios.
@@
In this paper, we primarily focus on multiple code-related tasks, aiming to generate high-quality and diverse instructional data tailored to specific task requirements. Addressing the aforementioned challenges, we refine the instruction data by classifying the instruction instances to four universal code-related tasks in CodeXGLUE Lu et al. (2021): 1) Code Summarization, 2) Code Generation, 3) Code Translation, 4) Code Repair and propose a widespread and versatile enhanced instruction generation method that could make full use of open source code data and stably generate high quality and diverse instruction data in multi-task scenarios. By this generation strategy, we obtain a dataset of 19,915 instruction instances across four code-related tasks, termed CodeSeaXDataset.
To validate our approach, we train StarCoder Li et al. (2023a), CodeLLaMa Roziere et al. (2023), and DeepseekCoder Guo et al. (2024) with our initial CodeSeaXDataset dataset and get WaveCoder. Following a thorough assessment on HumanEval Chen et al. (2021), MBPP Austin et al. (2021), HumanEvalPack Muennighoff et al. (2024) benchmarks, experimental results show that our WaveCoder exhibits outstanding generalization ability based on widespread and versatile enhanced instruction tuning. Moreover, to further explore the improvements brought by data quality, we use GPT-4 OpenAI (2023) to regenerate response for the instruction in CodeSeaXDataset. Fine-tuned with the enhanced 20K CodeSeaXDataset dataset, we obtain WaveCoder-Pro-6.7B which achieve 72.0% pass@1 on HumanEval Chen et al. (2021) and surpass open source Code LLMs but still behind SoTA Code LLM. Combining enhanced CodeSeaXDataset with WaveCoder-evol-codealpaca, the decontaminated Magicoder-evol-codealpaca 222https://huggingface.co/datasets/ise-uiuc/Magicoder-evol-codealpaca-110K dataset, we present WaveCoder-Ultra-6.7B, with SoTA generalization capabilities on multiple code-related tasks.
2 CodeSeaXDataset: Four-task Code-related Instruction Data
2.1 Tasks Details
| Task | Num | Per(%) | Prompt |
| Code Generation | 11370 | 57.1 | Implementing functions that perform specific operations given input. |
| Code Summarization | 3165 | 15.8 | Write clear and concise documentation for the given code. |
| Code Repair | 3144 | 15.8 | Identify and fix errors in the given code. |
@@ 2.1 Tasks Details
Given the code-related tasks from CodeXGLUE Lu et al. (2021), we select four of the most universally representative and common tasks from the three generative tasks (code-to-text, text-to-code, and code-to-code) for further exploration including Code Summarization, Code Generation, Code Translation, and Code Repair. Detailed descriptions of these tasks can be found below.
@@
Code Summarization (code-to-text). This task aims to create a brief summary of a given code. The raw code is used as input and the teacher model’s response is reformulated into an instruction format.
Code Generation (text-to-code, code-to-code). In this task, the model is expected to generate code based on a user’s demand description. Therefore, the teacher model is expected to generate instructions and solution code given the raw code as a instruction-solution pair. The generated solution code is then considered as the output.
Code Translation (code-to-code). This task involves converting one programming language into another. The task-specific prompt and raw code are given to the teacher model, then the model generates instructions and the translated code.
Code Repair (code-to-code). The aim of this task is to provide correct code based on potential issues in the given code. The teacher model is expected to generate solutions for the incorrect code, typically with the correct code and some descriptions, which are then taken as the output.
2.2 Widespread and Versatile Enhanced Instruction Generation
In past research work Zhou et al. (2023); Gupta et al. (2023), many researchers have discovered that data quality and diversity often play a more important role in instruction tuning process than data amount. The improvement of data quality and diversity are directly related to the performance of the fine-tuned LLM. Therefore, to ensure the data quality and diversity of instruction instance, we propose a widespread and versatile enhanced instruction generation method including two the following parts: 1) a method that can retain the diversity of instruction data by retainig the diversity of raw code to the utmost extent. 2) a LLM-based Generator-Discriminator framework to stably generate high-quality instruction data.
@@ 2.2.1 Raw Code Collection
To ensure the quality and diversity of raw code, we manually define some filtering rules and utilize a cluster method KCenterGreedy Sener and Savarese (2018); Chen et al. (2023) to get the raw code collection from the open source code dataset. In this work, we select CodeSearchNet 333https://huggingface.co/datasets/code_search_net, which contains 2 million of <comment, code> pairs from open-source libraries hosted on GitHub, as our foundation dataset and process it with the following steps:
Manually defined filtering rules. In order to select high-quality code for instruction-tuning, we make the following rules to filter the foundation dataset: i) In this work, we filtered the code to make sure that the length of required code is neither too long nor too short. ii) Followed Code Alpaca Chaudhary (2023), we have eliminated the raw code containing words from the blacklist, which could potentially reduce the performance of the resulting model.
Coreset selection method. To ensure the data diversity when select raw code samples, we employed KCenterGreedy Sener and Savarese (2018) algorithm, which has been proven efficient in obtaining a set of core samples of one distribution, to select representative samples from the open source code dataset based on the code embeddings encoded by the same embedding model (roberta-large-v1 Liu et al. (2019)).
By incorporating such a method into the open source code dataset, the diversity of the generated data no longer relies solely on capability of the teacher LLM itself or initial seed. Moreover, due to the application of the KCenterGreedy algorithm, the diversity of languages is also significantly retained, as shown in Table 2.
@@
| OctoCoder | 15B | StarCoder | 13K | 46.2 | 43.5 |
| WizardCoder | 15B | StarCoder | 78K | 57.3 | 51.8 |
| WaveCoder-SC-15B |
@@
| CodeLLaMa | 7B | - | ✘ | 33.5 | 41.4 |
| CodeLLaMa-instruct | 7B | CodeLLaMa | 14K | 34.8 | 44.4 |
| WaveCoder-CL-7B |
@@
| CodeLLaMa | 13B | - | ✘ | 36.0 | 47.0 |
| CodeLLaMa-instruct | 13B | CodeLLaMa | 14K | 42.5 | 49.4 |
| WaveCoder-CL-13B |
@@
| DeepseekCoder | 6.7B | - | ✘ | 49.4 | 60.6 |
| Magicoder-DS | 6.7B | DeepseekCoder | 75K | 66.5 | 60.4 |
| WaveCoder-DS-6.7B |
@@
| WaveCoder-Pro-6.7B |
6.7B | DeepseekCoder | 20K | 72.0 (+22.6) | 63.6 (+3.0) |
| SoTA Open-Source Models | |||||
| DeepseekCoder-instruct∗ | 6.7B | DeepseekCoder | - | 73.8 | 62.8 |
| Magicoder-S-DS | 6.7B | DeepseekCoder | 185K | 76.8 | 64.6 |
| WaveCoder-Ultra-6.7B |
6.7B | DeepseekCoder | 130K | 78.6 (+29.2) | 64.4 (+3.8) |
2.2.2 LLM-based Generator-Discriminator Framework
After the process of raw code collection, the data diversity from raw code has been retained, where the next step is to generate instruction data for supervised fine-tuning from the raw code. To further ensure the quality of generated instruction data, shown in Figure 2, we propose a LLM-based Generator-Discriminator framework where the generator can leverage an extensive amount of unsupervised open source code to generate supervised instruction data and the discriminator can generate analysis for each component in instruction data.
Generation Phase. In the generation phase, we utilize GPT-4 to generate definitions for each code-related task. As shown in Figure 2, following the model-generated task definition, we manually develop the generation requirements for the each code-related task. Integrating both the task definition and all associated requirements into the generation prompt, we take the raw code as an input and select different examples from the example database to generate instruction data by GPT-3.5.
Discrimination Phase. During the exploration of the instruction generation process, we noticed that the data quality of the instruction instances cannot be guaranteed through the generation phase alone. In order to enhance the controllability of data generation and further ensure the data quality, we employ GPT-4 as a LLM-based discriminator to continue analyzing and filtering the instruction data. Subsequently, inspired by Zero-shot-CoT Kojima et al. (2022), we establish a series of rules, exemplified in Figure 4 and disassemble them to some subtopics to ensure the discriminating accuracy where the LLM-based discriminator can analyze the generation step by step. By adopting this method, the discrimination rules can be modified partially to address certain issues. After the discrimination process, as shown in Figure 1, each instruction instance is classified as either a good or bad case and the classification information is subsequently random selected in the following generation as examples. For the reusage of these classified instruction instance, different from self-instruct Wang et al. (2023a) which solely utilize the initial seed task as good example, we exploit both the good generation and bad generation as few-shot example so that the generator can learn from the mistake in different bad example. Therefore, this framework provides a comprehensive approach to generating and evaluating instruction data, ensuring a high-quality training dataset.